SPECTROSCOPY AND DYNAMICS OF JET-COOLED P-COUMARIC ACID ESTERS: AN INVESTIGA-TION INTO THE PHOTOISOMERIZATION OF THE PHOTOACTIVE YELLOW PROTEIN

\underline{W. J. BUMA}, M. de GROOT, University of Amsterdam, Faculty of Science, Van ’t Hoff Institute for Molecular Sciences, Nieuwe Achtergracht 166, 1018 WV Amsterdam, The Netherlands.

Trans to cis photoisomerization of the chromophore of the photoactive yellow protein (PYP) is the first of a chain of events that leads to the negative phototaxis to blue light of certain purple bacteria. The spectroscopy and dynamics of jet-cooled p-coumaric acid methylester and the p-coumaric acid methythiolester, as well as of their clusters with water have been studied using mass-resolved REMPI and LIF experiments in combination with ab initio calculations. Because the esters more closely mimic the chromophore of PYP, we believe that these experiments contribute to a better understanding of the photoisomerization dynamics of PYP.